# LoRA-Pro: Are Low-Rank Adapters Properly Optimized?

### v1

@@ Abstract
Low-Rank Adaptation, also known as LoRA, has emerged as a prominent method for parameter-efficient fine-tuning foundation models by re-parameterizing the original matrix into the product of two low-rank matrices. Despite its efficiency, LoRA often yields inferior performance compared to full fine-tuning. In this paper, we propose LoRA-Pro to bridge this performance gap.
Firstly, we delve into the optimization processes in LoRA and full fine-tuning. We reveal that while LoRA employs low-rank approximation, it neglects to approximate the optimization process of full fine-tuning. To address this, we introduce a novel concept called the "equivalent gradient." This virtual gradient makes the optimization process on the re-parameterized matrix equivalent to LoRA, which can be used to quantify the differences between LoRA and full fine-tuning. The equivalent gradient is derived from the gradients of matrices $A$ and $B$. To narrow the performance gap, our approach minimizes the differences between the equivalent gradient and the gradient obtained from full fine-tuning during the optimization process. By solving this objective, we derive optimal closed-form solutions for updating matrices $A$ and $B$. Our method constrains the optimization process, shrinking the performance gap between LoRA and full fine-tuning.
Extensive experiments on natural language processing tasks validate the effectiveness of our method.
@@ 1 Introduction
Foundational models [Radford et al., 2021, Brown et al., 2020, Achiam et al., 2023, Kirillov et al., 2023, Rombach et al., 2022] have become the cornerstone of modern deep learning. By undergoing pre-training on massive datasets, these models typically exhibit excellent generalization and versatility. Remarkably, some foundation models even demonstrate emergent properties [Hoffmann et al., 2022, Kaplan et al., 2020]. As a result, foundation models have been widely applied to various downstream applications.
@@
Despite these advantages, the huge number of parameters in foundational models hinders their broader application. The substantial parameter count results in high fine-tuning costs for these tasks. To address this issue, recent research has focused on parameter-efficient fine-tuning (PEFT) methods [Hu et al., 2022, Houlsby et al., 2019, Lester et al., 2021, Zhou et al., 2022]. PEFT methods reduce the fine-tuning cost by keeping the foundation models frozen and only fine-tuning small, additional lightweight adapters. With the majority of parameters frozen, PEFT enables faster fine-tuning and requires fewer computational resources.
@@
Low-rank adaptation [Hu et al., 2022], also known as LoRA, is one of the most famous PEFT methods, which has been widely adopted across various domains. Inspired by previous works [Aghajanyan et al., 2021, Li et al., 2018], LoRA hypothesizes that the changes in weights during model adaptation exhibit a low-rank structure. To capture this, LoRA re-parameterizes these changes by expressing them as the product of two low-rank matrices: $W=W_{0}+\Delta W\approx W_{0}+sBA$, where $s$ is a scaling factor, and $A\in\mathbb{R}^{r\times n}$ and $B\in\mathbb{R}^{m\times r}$ are low-rank matrices with rank $r\ll\min(m,n)$. LoRA reduces the number of trainable parameters from $m\times n$ to $r\times(m+n)$, thereby decreasing the cost of fine-tuning. However, despite its efficiency, LoRA’s fine-tuning performance often falls short compared to full fine-tuning [Hu et al., 2022, Liu et al., 2024, Ding et al., 2023].
In this paper, we propose a novel PEFT method, LoRA-Pro, aimed at bridging the gap between LoRA and full fine-tuning. While LoRA employs low-rank approximation by re-parametrizing weight changes as the product of two low-rank matrices, it falls short in approximating the optimization process of full fine-tuning. To measure their discrepancy in the optimization process, we propose a novel concept, “Equivalent Gradient", for LoRA optimization. Equivalent gradient characterizes the gradient of the original matrix after low-rank approximation (despite it not being directly trainable), is composed of gradients from matrices A and B. Thus, during LoRA fine-tuning, our goal is not only to approximate the matrix with low-rank matrices but also to minimize the difference between the equivalent gradient and the gradient from full fine-tuning during the gradient descent process. This is achieved by selecting appropriate gradients for matrices A and B, ensuring a more accurate and effective fine-tuning process. To achieve this, we formulate it as an optimization problem. We then derive theoretical solutions for the problem, presenting optimal gradients for updating matrices A and B. These solutions ensure that the equivalent gradient closely match the optimization dynamics of full fine-tuning. By doing so, we enhance the effectiveness LoRA, bridging the gap between LoRA and full fine-tuning.
@@
  • We identify that LoRA approximates low-rank matrices but neglects to approximate the optimization process of full parameter fine-tuning. This shortcoming is one of the reasons for the performance gap between LoRA and full fine-tuning.
  • We introduce the concept of Equivalent Gradient, which allows us to quantify the discrepancy in the optimization process between LoRA and full fine-tuning. By minimizing this discrepancy, we derive the optimal closed-form updated solutions for LoRA.
  • Extensive experiments on natural language processing tasks validate the effectiveness of our method.
2 Related Work
Parameter-Efficient Fine-Tuning. Given the huge size of foundation models, recent research has focused on developing parameter-efficient fine-tuning methods [Hu et al., 2022, Liu et al., 2024, Ding et al., 2023, Houlsby et al., 2019, Liu et al., 2023, Lester et al., 2021]. These methods aim to reduce the cost of fine-tuning by adjusting only a small portion of the model’s parameters. Generally, these methods fall into two main categories. The first category is adapter tuning [Houlsby et al., 2019, Sung et al., 2022, He et al., 2021, Zhang et al., 2024, Bapna and Firat, 2019, Hu et al., 2022], which involves inserting small neural network modules, called adapters, into specific layers of the model. During fine-tuning, we keep the model frozen and only fine-tune the lightweight adapter modules, significantly reducing the memory footprint for fine-tuning. The second category is prompt tuning [Lester et al., 2021, Zhou et al., 2022, Li and Liang, 2021, Liu et al., 2022]. Prompt tuning adapts the models to specific tasks by adding specially designed prompts or learnable tokens to the input data, rather than directly modifying the internal parameters of foundation models. In this paper, we focus on LoRA [Hu et al., 2022], a prominent method within the realm of adapter tuning.
Low Rank Adaptation. Low-rank adaptation, initially referred to as LoRA [Hu et al., 2022], has evolved into a broad category encompassing parameter-efficient fine-tuning methods based on low-rank approximations [Hu et al., 2022, Liu et al., 2024, Hayou et al., 2024, Kalajdzievski, 2023, Zhang et al., 2023, Kopiczko et al., 2024, Hyeon-Woo et al., 2022, Zhang and Pilanci, 2024, Wang et al., 2024, Zhao et al., 2024]. LoRA [Hu et al., 2022] assumes that the changes in the weights of pre-trained models exhibit a low-rank structure. Consequently, it re-parameterizes these changes as the product of low-rank matrices, thereby reducing the cost associated with fine-tuning.
Several variants of LoRA have been proposed to address different aspects of this approach. For example, DoRA [Liu et al., 2024] improves LoRA [Hu et al., 2022] by incorporating a learnable magnitude vector to re-scale the normalized product of low-rank matrices. Another variant, rsLoRA Kalajdzievski [2023], introduces a new scaling factor to stabilize training in high-rank scenarios. LoRA+[Hayou et al., 2024] improves upon LoRA by applying different learning rates to the two low-rank matrices. Additionally, Galore [Zhao et al., 2024] employs SVD to project the gradients of full parameter training into a low-rank space, thereby reducing the memory footprint during pre-training and fine-tuning.
@@ 3 Method
In this section, we begin by revisiting LoRA [Hu et al., 2022] in Section 3.1. Following this, we conduct a comparison between LoRA and full fine-tuning from an optimization perspective in Section 3.2. Finally, in Section 3.3, we point out that LoRA falls short in approximating full fine-tuning during the optimization process, and we introduce LoRA-Pro as a solution to bridge this performance gap.
3.1 Revisit Low Rank Adaptation
First of all, let’s dive back into Low-Rank Adaptation (LoRA) [Hu et al., 2022]. LoRA’s core idea revolves around recognizing the low-rank structure of the change matrix $\Delta W$ in the standard fine-tuning process. This insight allows LoRA [Hu et al., 2022] to re-parameterize the change matrix into the product of two low-rank matrices,
$W=W_{0}+\Delta W=W_{0}+sBA.$ (1)
Here, $W_{0}\in\mathbb{R}^{m\times n}$ represents the pre-trained weight matrix, $B\in\mathbb{R}^{m\times r}$ and $A\in\mathbb{R}^{r\times n}$ are the low-rank matrices, and $s$ is a scaling factor. For LoRA [Hu et al., 2022], $s=\frac{\alpha}{r}$, while for rsLoRA [Kalajdzievski, 2023], $s=\frac{\alpha}{\sqrt{r}}$. Here, $\alpha$ is the hyper-parameter and $r\ll min(m,n)$ denotes the rank. Consequently, LoRA significantly reduces the number of fine-tuning parameters from $m\times n$ to $r\times(m+n)$.
@@ 3.1 Revisit Low Rank Adaptation
Despite widespread applications across various domains, LoRA’s performance still falls short when compared to full fine-tuning. In this part, we review and compare LoRA and full fine-tuning in the optimization process. In full fine-tuning, we utilize differential to analyze the relationship between changes in the loss and changes in the weights:
@@
[…]
In LoRA optimization, given that $W=W_{0}+sBA$, we compute the differential using the chain rule:
@@
[…]
@@ 3.3 Low-Rank Adaptation with Equivalent Gradient
[…]
In this section, Equivalent Gradient. From Equation (3), we can see that changes in matrices $A$ and $B$ are inherently linked to changes in matrix $W$ through the chain rule:
$\mathrm{d}W=\frac{\partial W}{\partial A}^{T}\mathrm{d}A+\frac{\partial W}{\partial B}^{T}\mathrm{d}B=-(sBg^{A}_{lora}+sg^{B}_{lora}A).$ (6)
In comparison to full fine-tuning, this is equivalent to updating $W$ using the gradient $\tilde{g}=sBg^{A}_{lora}+sg^{B}_{lora}A$. This critical relationship has been neglected in the LoRA optimization process. Hence, we hypothesize that by carefully adjusting the gradients of matrices $A$ and $B$ in such a way that $\tilde{g}$ under LoRA closely approximates the gradient $g$ from full fine-tuning, we can effectively bridge the gap between LoRA and full fine-tuning.
Based on this relationship, we define the concept of equivalent gradient in Definition 3.3. Equivalent gradient describes the gradient of the matrix $W$ following low-rank adaptation, despite $W$ not being a trainable parameter. To narrow the performance gap, our goal is to carefully select suitable $g^{A}$ and $g^{B}$ to minimize the distance between the equivalent gradient $\tilde{g}$ and the gradient under full fine-tuning $g$. Hence, our objective is:
@@
Theorem 4.1 Assume matrices $B\in\mathbb{R}^{m\times r},A\in\mathbb{R}^{r\times n}$ are both full rank. For the objective $\min_{g^{A},g^{B}}\|\tilde{g}-g\|^{2}_{F}$, the solutions are given by: $g^{A}$ $=\frac{1}{s}(B^{T}B)^{-1}B^{T}g+XA=\frac{1}{s^{2}}(B^{T}B)^{-1}g^{A}_{lora}+XA$ (8) $g^{B}$ $=\frac{1}{s}[I-B(B^{T}B)^{-1}B^{T}]gA^{T}(AA^{T})^{-1}-BX=\frac{1}{s^{2}}[I-B(B^{T}B)^{-1}B^{T}]g^{B}_{lora}(AA^{T})^{-1}-BX.$ (9) Here, $X\in\mathbb{R}^{r\times r}$ represents an arbitrary matrix.
Closed-form Solution. Fortunately, Equation (7) admits a closed-form solution. According to Theorem 3.3, we obtain the optimal gradients for matrices $A$ and $B$, ensuring that the equivalent gradient achieves the best approximation to the full fine-tuning gradient. Moreover, we observe that $g^{A}$ and $g^{B}$ can be expressed as $g^{A}_{lora}$ and $g^{B}_{lora}$, respectively, indicating that we do not explicitly possess the full fine-tuning gradient $g$. Therefore, our approach involves back-propagating in standard LoRA and adjusting the gradients of matrices $A$ and $B$ using the closed-form solution outlined in Theorem 3.3.
Theorem 4.2 When updating matrices $A$ and $B$ using the closed-form solution from Theorem 3.3, we proceed as follows: $A\leftarrow A-\gamma g^{A}$ (10) $B\leftarrow B-\gamma g^{B},$ (11) where $\gamma\geq 0$ denotes the learning rate. Our method ensures a decrease in the loss, akin to the standard gradient descent algorithm, expressed by: $\mathrm{d}L=-\gamma\{\langle g^{A}_{lora},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{A}_{lora}\rangle_{F}+\langle g^{B}_{lora},\frac{1}{s^{2}}[I-B(B^{T}B)^{-1}B^{T}]g^{B}_{lora}(AA^{T})^{-1}\rangle_{F}\}\leq 0$ (12)
@@
Although Theorem 3.3 provides a closed-form solution to the optimization problem $\min_{g^{A},g^{B}}\|\tilde{g}-g\|^{2}_{F}$, this does not necessarily mean that updating matrices $A$ and $B$ with this solution will decrease the loss. To address this, we have Theorem 3.3, which guarantees a decrease in the loss during the optimization process. This theorem indicates that the change in loss, $\mathrm{d}L$, can be expressed as a negative scalar multiplied by the sum of two positive definite quadratic forms. This relationship ensures that $\mathrm{d}L\leq 0$ during the update process, thus consistently driving the optimization process towards a lower loss.
Theorem 4.3 Consider the optimization problem, $\min_{X}\|g^{A}-g^{A}_{lora}\|_{F}^{2}+\|g^{B}-g^{B}_{lora}\|_{F}^{2},$ (13) where $g^{A}$ and $g^{B}$ are the optimal solutions as stated in Theorem 3.3. The optimal $X$ can be determined by solving the Sylvester equation: $B^{T}BX+XAA^{T}=-\frac{1}{s^{2}}(B^{T}B)^{-1}g^{A}_{lora}A^{T},$ (14) which has a unique solution $X$ provided that $B^{T}B$ and $-AA^{T}$ do not have any shared eigenvalues.
@@
Selection of X. Although the equivalent gradient itself is not directly related to the matrix $X$, the presence of $X$ plays a significant role in the updates of matrices $A$ and $B$. We select an appropriate $X$ such that $g^{A}$ and $g^{B}$ remain close to $g^{A}_{lora}$ and $g^{B}_{lora}$ respectively. Consequently, we minimize their Frobenius norm, as demonstrated in Equation (41). In practical terms, $B^{T}B$ and $AA^{T}$ do not share common eigenvalues. Therefore, according to Theorem 3.3, we can determine a unique optimal $X$ for updating matrices $A$ and $B$.

### v2

@@ Abstract
Low-rank adaptation, also known as LoRA, has emerged as a prominent method for parameter-efficient fine-tuning of foundation models. Despite its computational efficiency, LoRA still yields inferior performance compared to full fine-tuning. In this paper, we first uncover a fundamental connection between the optimization processes of LoRA and full fine-tuning: using LoRA for optimization is mathematically equivalent to full fine-tuning using a low-rank gradient for parameter updates. And this low-rank gradient can be expressed in terms of the gradients of the two low-rank matrices in LoRA. Leveraging this insight, we introduce LoRA-Pro, a method that enhances LoRA’s performance by strategically adjusting the gradients of these low-rank matrices. This adjustment allows the low-rank gradient to more accurately approximate the full fine-tuning gradient, thereby narrowing the performance gap between LoRA and full fine-tuning. Furthermore, we theoretically derive the optimal solutions for adjusting the gradients of the low-rank matrices, applying them during fine-tuning in LoRA-Pro. We conduct extensive experiments across natural language understanding, dialogue generation, mathematical reasoning, code generation, and image classification tasks, demonstrating that LoRA-Pro substantially improves LoRA’s performance, effectively narrowing the gap with full fine-tuning. Code is publicly available at https://github.com/mrflogs/LoRA-Pro.
@@ 1 Introduction
Foundational models (Radford et al., 2021; Brown et al., 2020; Achiam et al., 2023; Kirillov et al., 2023; Rombach et al., 2022; Touvron et al., 2023) have become the cornerstone of modern deep learning. By undergoing pre-training on massive datasets, these models typically exhibit excellent generalization and versatility. Remarkably, some foundation models even demonstrate emergent properties (Hoffmann et al., 2022; Kaplan et al., 2020). Due to these advantages, foundational models have been widely applied to various downstream applications.
@@
Nevertheless, it still requires additional fine-tuning when applied to downstream tasks, where the huge parameter size of foundation models result in high cost in this stage. To address this issue, recent research has focused on parameter-efficient fine-tuning (PEFT) methods (Hu et al., 2022; Houlsby et al., 2019; Lester et al., 2021). PEFT methods reduce the fine-tuning cost by keeping the foundation models frozen and only fine-tuning small, additional lightweight adapters. With the majority of parameters frozen, PEFT enables faster fine-tuning and requires fewer resources.
@@
Low-rank adaptation (Hu et al., 2022), also known as LoRA, is one of the most famous PEFT methods, which has been widely adopted across various domains. Inspired by previous works (Aghajanyan et al., 2021; Li et al., 2018), LoRA hypothesizes that the changes in weights during model adaptation exhibit a low-rank structure. To capture this, LoRA re-parameterizes these changes by expressing them as the product of two low-rank matrices: $W=W_{0}+\Delta W\approx W_{0}+sBA$, where $s$ is a scaling factor, and $A\in\mathbb{R}^{r\times n}$ and $B\in\mathbb{R}^{m\times r}$ are low-rank matrices with rank $r\ll\min(m,n)$. LoRA reduces the number of trainable parameters from $m\times n$ to $r\times(m+n)$, thereby decreasing the cost of fine-tuning. However, despite its efficiency, LoRA’s fine-tuning performance often falls short compared to full fine-tuning (Hu et al., 2022; Liu et al., 2024; Ding et al., 2023).
In this paper, we propose a novel PEFT method, LoRA-Pro, aimed at bridging the gap between LoRA and full fine-tuning. To begin with, we uncover a crucial connection between the optimization processes of LoRA and full fine-tuning: using LoRA for optimization is equivalent to full fine-tuning using a low-rank gradient for parameter updates. In LoRA, we discover that the change in weight $W$ is connected to the changes in matrices $A$ and $B$, expressed as $\mathrm{d}W=\frac{\partial W}{\partial A}\mathrm{d}A+\frac{\partial W}{\partial B}\mathrm{d}B$. This relationship implies that updating matrices $A$ and $B$ with gradients $g^{A}$ and $g^{B}$ is equivalent to updating $W$ with a low-rank equivalent gradient $\tilde{g}$ in full fine-tuning, where:
$\tilde{g}=\frac{\partial W}{\partial A}g^{A}+\frac{\partial W}{\partial B}g^{B}=sBg^{A}+sg^{B}A.$ (1)
Leveraging this insight, our goal is to bridge LoRA’s gap with full fine-tuning by minimizing the discrepancy between the low-rank equivalent gradient $\tilde{g}$ and the full fine-tuning gradient $g$, by adjusting the gradients of matrices $A$ and $B$, i.e., $\min_{g^{A},g^{B}}\|\tilde{g}-g\|^{2}_{F}$. Furthermore, we theoretically demonstrate that this optimization problem admits an optimal closed-form solution, as shown in Theorem 2.1. Notably, the optimal gradients for the low-rank matrices do not explicitly depend on the full fine-tuning gradient.
@@
  • We first uncover a crucial connection between LoRA and full fine-tuning in the optimization process: optimizing with LoRA is mathematically equivalent to full fine-tuning using a low-rank gradient for updating.
  • We propose a novel PEFT method called LoRA-Pro. Our approach minimizes the distance between the true gradient and the low-rank gradient by adjusting the gradients of matrices A and B. We theoretically prove the optimal gradients and optimize using these gradients.
  • Extensive experiments across tasks in natural language understanding, dialogue generation, mathematical reasoning, code generation, and image classification demonstrate the effectiveness of our method.
@@ 2 Method
In this section, we begin by revisiting LoRA (Hu et al., 2022) in Section 2.1. Following this, we conduct a comparison between LoRA and full fine-tuning, and point out their connection in the optimization process in Section 2.2. Finally, in Section 2.3, we introduce LoRA-Pro as a solution to bridge the gap between LoRA and full fine-tuning.
2.1 Revisiting Low-Rank Adaptation
First of all, let’s dive back into Low-Rank Adaptation (LoRA) (Hu et al., 2022). LoRA’s core idea revolves around recognizing the low-rank structure of the change matrix $\Delta W$ in the standard fine-tuning process. This insight allows LoRA (Hu et al., 2022) to re-parameterize the change matrix into the product of two low-rank matrices,
$W=W_{0}+\Delta W\approx W_{0}+sBA.$ (2)
Here, $W_{0}\in\mathbb{R}^{m\times n}$ represents the pre-trained weight matrix, $B\in\mathbb{R}^{m\times r}$ and $A\in\mathbb{R}^{r\times n}$ are the low-rank matrices, and $s$ is a scaling factor. For LoRA (Hu et al., 2022), $s=\frac{\alpha}{r}$, while for rsLoRA (Kalajdzievski, 2023), $s=\frac{\alpha}{\sqrt{r}}$. Here, $\alpha$ is the hyper-parameter and $r\ll min(m,n)$ denotes the rank. Consequently, LoRA significantly reduces the number of fine-tuning parameters from $m\times n$ to $r\times(m+n)$, thereby decreasing the computational cost of fine-tuning.
@@ 2.1 Revisiting Low-Rank Adaptation
Despite its widespread applications across various domains, LoRA’s performance still falls short when compared to full fine-tuning. In this part, we compare LoRA and full fine-tuning in the optimization process. Then, we demonstrate that optimizing using LoRA is equivalent to using a low-rank gradient in full fine-tuning for updating the parameters.
Full fine-tuning. In full fine-tuning, we utilize differential to analyze the relationship between changes in the loss and changes in the weights:
@@
where $\mathrm{d}L$ and $\mathrm{d}W$ denotes the changes of the parameter $W$ and the loss $L$, and $\langle\cdot,\cdot\rangle_{F}$ is the Frobenius inner product. To minimize the loss function, we typically set $\mathrm{d}W=-\frac{\partial L}{\partial W}\triangleq-g$ (omitting the learning rate for simplicity), which results in $\mathrm{d}L=-\|\frac{\partial L}{\partial W}\|^{2}_{F}\leq 0$.
Low-rank adaptation. In LoRA optimization, given that $W=W_{0}+sBA$, we compute the differential using the chain rule:
@@
Similarly, LoRA sets $\mathrm{d}A=-\frac{\partial L}{\partial A}\triangleq-g^{A}_{lora}$ and $\mathrm{d}B=-\frac{\partial L}{\partial B}\triangleq-g^{B}_{lora}$, and thus $\mathrm{d}L=-\|\frac{\partial L}{\partial A}\|^{2}_{F}-\|\frac{\partial L}{\partial B}\|^{2}_{F}\leq 0$. Moreover, employing the chain rule, we derive:
@@
Why LoRA performs worse than full fine-tuning. With Equation (3) and (4), we observe a critical connection between full fine-tuning and LoRA in the optimization process. In LoRA, changes in matrices $A$ and $B$ are inherently linked to changes in matrix $W$, i.e., $\mathrm{d}W=\frac{\partial W}{\partial A}^{T}\mathrm{d}A+\frac{\partial W}{\partial B}^{T}\mathrm{d}B$. This indicates that updating $A$ and $B$ with gradient $g^{A}$ and $g^{B}$ is equivalent to performing full fine-tuning on $W$ via the following update:
$\mathrm{d}W=\frac{\partial W}{\partial A}^{T}\mathrm{d}A+\frac{\partial W}{\partial B}^{T}\mathrm{d}B=-(sBg^{A}+sg^{B}A).$ (6)
Equation (6) reveals that LoRA optimization is equivalent to full fine-tuning using a low-rank gradient $\tilde{g}=sBg^{A}+sg^{B}A$ (which rank is at most to $2r$ [^1] We provide the proof in Appendix B.1) for optimization. Consequently, the performance gap between LoRA and full fine-tuning may stem from differences between $\tilde{g}$ and the full gradient $g$. The low-rank gradient $\tilde{g}$ may lose important information contained in $g$, leading to distinct optimization trajectories and ultimately causing LoRA to converge to a sub-optimal solution.
@@ 2.3 Low-Rank Adaptation with Equivalent Gradient
Definition 1 (Equivalent Gradient).
In the context of LoRA optimization, we define the equivalent gradient as, $\tilde{g}\triangleq sBg^{A}+sg^{B}A,$ where $s$ is the scaling factor, and $g^{A}$ and $g^{B}$ are gradients with respect to $A$ and $B$, respectively.
In this part, we introduce our LoRA-Pro method, which bridges the performance gap by minimizing the discrepancy between the gradients. For convenience, we define the concept of equivalent gradient in Definition 1. Equivalent gradient describes the virtual low-rank gradient of the matrix $W$ in LoRA optimization process, despite $W$ not being a trainable parameter. To narrow the performance gap, our goal is to carefully adjust $g^{A}$ and $g^{B}$ of matrices $A$ and $B$ to minimize the distance between the equivalent gradient $\tilde{g}$ and the full gradient $g$ in full fine-tuning. Hence, our objective is:
@@
Here, $\|\cdot\|_{F}$ denotes the Frobenius norm, and $\mathrm{d}L$ denotes the change in loss when updating with gradients $g^{A}$ and $g^{B}$. The objective aims to minimize the distance of the gradients while ensuring a decrease in loss using the solutions for $g^{A}$ and $g^{B}$.
Figure 1: Illustration of LoRA-Pro. LoRA (Hu et al., 2022) reduces the trainable parameter by re-parameterizing the weight into the product of two low-rank matrices, i.e., $W=W_{0}+sBA$. We have discovered a connection between the optimization processes of full fine-tuning and LoRA. Updating matrices $B$ and $A$ using gradients $g^{B}$ and $g^{A}$ is equivalent to updating weight $W$ using a virtual low-rank gradient $\tilde{g}=sBg^{A}+sg^{B}A$. Therefore, in LoRA-Pro, we aim to adjust gradients $g^{B}$ and $g^{A}$ to minimize the distance between the equivalent gradient $\tilde{g}$ and the full fine-tuning gradient $g$, thereby reducing their performance gap. In Theorem 2.1, we provide the optimal update gradients, and in Appendix C, we present the pseudo-code for the optimization algorithm.
Closed-form solution. Fortunately, we prove that the optimization problem (7) admits an optimal closed-form solution, as stated in Theorem 2.1. Additionally, an interesting observation arises from Theorem 2.1: while the full gradient $g$ serves as the ground truth in the objective, it does not necessarily explicit appear in the closed-form solution. Instead, the closed-form solution for the optimal gradients can be expressed in terms of the gradients of LoRA. This allows for an efficient gradient adjustment process, where we backpropagate using the standard LoRA and adjust the gradients of matrices $A$ and $B$ based on the closed-form solution presented in Theorem 2.1.  [^2] We provide detailed algorithms in Appendix C.
Theorem 2.1.
@@
Assume matrices $B\in\mathbb{R}^{m\times r},A\in\mathbb{R}^{r\times n}$ are both full rank. For the objective $\min_{g^{A},g^{B}}\|\tilde{g}-g\|^{2}_{F}$, the optimal solutions are given by: $g^{A}$ $=\frac{1}{s}(B^{T}B)^{-1}B^{T}g+XA=\frac{1}{s^{2}}(B^{T}B)^{-1}g^{A}_{lora}+XA,$ (8) $g^{B}$ $=\frac{1}{s}[I-B(B^{T}B)^{-1}B^{T}]gA^{T}(AA^{T})^{-1}-BX$ (9) $=\frac{1}{s^{2}}[I-B(B^{T}B)^{-1}B^{T}]g^{B}_{lora}(AA^{T})^{-1}-BX.$ (10) Here, $X\in\mathbb{R}^{r\times r}$ represents an arbitrary matrix.
Proof.
See Appendix B.2. ∎
@@
Loss minimization. While Theorem 2.1 offers a closed-form solution to the optimization problem $\min_{g^{A},g^{B}}\|\tilde{g}-g\|^{2}_{F}$, it’s crucial to understand that this solution does not inherently guarantee a decrease in loss when updating the matrices $A$ and $B$. To address this concern, we introduce Theorem 2.2. In this theorem, we prove that the change in loss $\mathrm{d}L$ can be expressed as a negative sum of two Frobenius norms, which leads to $\mathrm{d}L<0$. This result ensures that the optimization process consistently drives towards a lower loss.
@@
Selection of matrix X. Although the equivalent gradient itself is not directly related to the matrix $X$, the presence of $X$ plays a significant role in the updates of matrices $A$ and $B$. We select an appropriate $X$ such that $g^{A}$ and $g^{B}$ remain close to $g^{A}_{lora}$ and $g^{B}_{lora}$ respectively. To achieve this, we minimize their Frobenius norm, as demonstrated in Equation (14). In practical terms, $B^{T}B$ and $-AA^{T}$ do not share common eigenvalues. Therefore, according to Theorem 2.3, we can determine a unique optimal $X$ for updating matrices $A$ and $B$.
Theorem 2.2.
When updating matrices $A$ and $B$ using the closed-form solution from Theorem 2.1, we proceed as follows: $A\leftarrow A-\gamma g^{A}$ (11) $B\leftarrow B-\gamma g^{B},$ (12) where $\gamma\geq 0$ denotes the learning rate. Our method ensures a decrease in the loss, akin to the standard gradient descent algorithm, expressed by: $\mathrm{d}L=-\gamma\{\langle g^{A}_{lora},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{A}_{lora}\rangle_{F}+\langle g^{B}_{lora},\frac{1}{s^{2}}[I-B(B^{T}B)^{-1}B^{T}]g^{B}_{lora}(AA^{T})^{-1}\rangle_{F}\}\leq 0.$ (13)
Proof.
See Appendix B.3. ∎
Theorem 2.3.
Consider the optimization problem, $\min_{X}\|g^{A}-g^{A}_{lora}\|_{F}^{2}+\|g^{B}-g^{B}_{lora}\|_{F}^{2},$ (14) where $g^{A}$ and $g^{B}$ are the optimal solutions as stated in Theorem 2.1. The optimal $X$ can be determined by solving the Sylvester equation: $B^{T}BX+XAA^{T}=-\frac{1}{s^{2}}(B^{T}B)^{-1}g^{A}_{lora}A^{T},$ (15) which has a unique solution $X$ provided that $B^{T}B$ and $-AA^{T}$ do not have any shared eigenvalues.
Proof.
See Appendix B.4. ∎